On January 1, 2020, Smith Corporation had the following balances in the balances in its’ stockholders’ equity

accounts.

Common Stock ($10 par value, 120,000 shares issued and outstanding) $1,200,000

Paid-in Capital in Excess of Par-Common Stock 200,000

Retained Earnings 600,000

The following transactions occurred during the year.

Jan. 1 Declared a $1 cash dividend per share to stockholders of record on January 15, payable February 1.

Feb. 1 Paid the dividend declared in January.

Mar. 1 Announced a 2-for-1 stock split. Prior to the split, the market price per share was $36.

June 1 Declared a 10% stock dividend to stockholders of record on June 15, distributable June 30. On June 1 the market price of the stock was $13 per share.

June 30 Issued the shares for the stock dividend.

Dec. 1 Declared a $.50 per share dividend to stockholders of record on December 15, payable January 10, 2016.

Dec 31 Determined that net income for the year was $350,000.

Instructions

a.Journalize the above transactions.

b. Journalize the closing entries for net income and dividends.

c.Enter the beginning balances in the stockholders’ equity accounts and post the above journal entries.

(You will need to open additional stockholders’ equity accounts as needed.)

d. Prepare the stockholders’ equity section of the balance sheet as of December 31.
Answer #1
Date Account Titles Debit Credit
Jan-01 Cash Dividend $              120,000 =120000*1
      Cash Dividend Payable $                120,000
(Cash dividend declared)
Feb-01 Cash Dividend Payable $              120,000
     Cash $                120,000
(Dividend paid in cash)
Mar-01 Common Stock ($10 par value, 120,000 shares issued and outstanding) $          1,200,000
Common Stock ($5 par value, 240,000 shares issued and outstanding) $            1,200,000
(Stock split 2 for 1)
Jun-01 Stock Dividend $              312,000 =240000*10%*13
      Stock Dividend Distributable $                120,000 =240000*10%*5
      Paid-in Capital in Excess of Par—Common Stock $                192,000 =312000-120000
(10% Common Stock Dividend declared)
Jun-30 Stock Dividend Distributable $              120,000
        Common Stock - $10 par value $                120,000
(Issued common stock for stock dividend)
Dec-01 Cash Dividend $              132,000 =(240000+24000)*0.5
      Cash Dividend Payable $                132,000
(Cash dividend declared)
Dec-31 Income Summary $              350,000
      Retained Earnings $                350,000
(Transferred net income to retained earnings)
Dec-31 Retained Earnings $              564,000
      Cash Dividend $                252,000
      Stock Dividend $                312,000
(To close dividend accounts)
Stockholder's Equity
Common Stock ($5 par value, 264,000 shares issued and outstanding) $          1,320,000
Paid in Capital in excess of par - Common stock $              392,000
Total Paid in capital $          1,712,000
Retained Earnings $              386,000
Total Stockholder's Equity $          2,098,000